how long must a rectangle be in order to hold 1800 cubic centimeters if its width is 12 centimeters and its height is 6 centimeters?

Answer by mananth(16946)

how long must a rectangle be in order to hold 1800 cubic centimeters if its width is 12 centimeters and its height is 6 centimeters?

volume = L * W * H
1800 = L * 12*6
1800/12*6 = L
L= 25 cm